the slope of a line is a -2/t and the line contains the points (7.4) and (x,12) what is the value of t

Answer by Alan3354(69443)

the slope of a line is a -2/t and the line contains the points (7.4) and (x,12) what is the value of t
----------
m = (12-4)/(x-7) = -2/t
8t = -2x + 14
t = -x/4 + 7/4
t = (7-x)/4